1. What is the mass in mg of 1.60 mmol of acetanilide? (Show your work)
Answer #1

Answer:

​​​​​​We'll use formula-

~ no. of moles =

or rearranging this we get,

mass = no. of moles × molar mass

1) calculating no. of moles -

we have 1.60 mmoles of acetanilide

1 mole = 1000 mmols

1mmol =

1.60mmol =

= 0.0016 moles

2) Calculating molar mass-

chemical formula of acetanilide = C8H9NO

So, molar mass = (12×8) + (1×9) + (14) + (16)

= 135 g/moles

3) Applying formula:

mass = molar mass × no. of moles

= (135 × 0.0016) grams

= 0.216 grams

or 1g = 1000 mg

0.216g = (0.216 × 1000) mg

​​​= 216 mg